# Olga Zatsarynna1, Emad Bahrami1, Yazan Abu Farha2, Gianpiero Francesca 3, Juergen Gall1, 4

### v1

@@ 4.3 Protocol and Evaluation
Table 1: Comparison of MANTA to the state-of-the-art methods on Breakfast, Assembly101 and 50Salads.
@@ 4.4 Results
Comparison to the State of the Art. In this section, we present the results of our proposed model, comparing its performance with the previous methods, as shown in Tab. 1. Our model achieves state-of-the-art results across all three datasets for both reported metrics. Specifically, on the Breakfast dataset, our method significantly outperforms previous approaches across all observation and anticipation horizons. On the Assembly101 dataset, MANTA surpasses all prior methods, with the exception of Top-1 MoC accuracy for $\beta=0.1$, where it performs on par with GTDA. Lastly, on the 50Salads dataset, our approach sets a new state-of-the-art for Mean MoC accuracy and Top-1 MoC accuracy for $\alpha=0.3$. For $\alpha=0.2$, our model’s Top-1 MoC is on par with earlier methods.
Beyond its effectiveness, our proposed model demonstrates substantial efficiency improvements compared to [55]. Fig. 4 compares the inference and training time of MANTA and the previous best-performing method GTDA [55] on the Breakfast dataset. On the left, it shows the average inference time required to generate $S=25$ samples with observation and anticipation ratios set to $\alpha=0.3$ and $\beta=0.5$ for videos in the test splits. While GTDA requires an average of $71.8$ seconds per video, our model needs only $1.1$ seconds, achieving a 65.3×65.3$\times$65.3 × speedup. Besides faster inference, our model significantly reduces training time as well, as shown in Fig. 4 on the right. The average runtime for a single epoch of GTDA is 379 seconds, while MANTA only takes 57 seconds per epoch. Additionally, due to the reduced number of the required blocks, our proposed model has 2.8×2.8$\times$2.8 × fewer parameters compared to GTDA, as shown in Tab. 6 and is therefore much more memory-efficient.
Finally, we demonstrate a qualitative comparison of our model to GTDA in Fig. 3. While GTDA anticipates future actions wrong, our proposed MANTA model predicts plausible action sequences. This demonstrates the limitations of the previous generator compared to our model, caused by its sparse receptive field lacking global context, leading not only to faulty predictions but also to over-segmentation.
Figure 3: Qualitative comparison on MANTA and GTDA on Breakfast. Best viewed zoomed in.
@@ 4.5 Ablations
Firstly, the selectivity of the BSS layer is enabled by employing the S6 layer for temporal modelling, which applies input-specific SSM matrices for data-driven sequence processing. To assess the impact of selectivity, we replaced the S6 layer with S4, for which the parameters remain fixed across timesteps. The results in Tab. 2 show a significant performance drop without selectivity (Ours w/o Sel.). This is expected, as the input sequences to our model consist of two distinct segments: observed and masked future frames. Processing these segments homogeneously is inherently sub-optimal, since it is not possible to adjust the contribution of individual values regardless of their validity and/or importance. GTDA made similar observations, where selectivity was ensured by data-driven temporal gating.
@@
MANTA Block. We further performed a series of experiments to explore how changes to the MANTA Block beyond the BSSL affect the performance of the final model. Firstly, we investigated the effect of the internal structure of the block, as shown in Tab. 4. We experimented with introducing a dilated convolution layer with an exponentially growing dilation rate (Ours w/ Dil. Conv) before the LN layer, akin to [55]. This, however, led to the degradation of results. Moreover, following recent works [17, 44], we experimented with modifying the channel mixing (FF) component of our block. Removing the mixing MLP altogether (Ours w/o FF) hinders the performance, as expected, due to a lack of channel-wise communication. However, introducing channel attention (Ours w/ SE), as proposed in [17], fails to improve the results and leads to a performance drop. Finally, we experimented with varying the total number of blocks in the final model (Tab. 5). Empirically, we found that the model with $B=15$ blocks showed the best results, with further increase or decrease in the number of blocks harming the model’s performance.
MoC Num. $\beta\ (\alpha=0.2)$ $\beta\ (\alpha=0.3)$
blocks 0.1 0.2 0.3 0.5 0.1 0.2 0.3 0.5
Mean 10 26.6 24.3 23.5 23.1 32.7 29.4 28.1 26.8
15 27.7 25.3 24.6 23.8 34.2 30.9 29.1 27.7
20 27.2 24.5 23.4 23.2 32.1 29.3 27.7 26.6
Top-1 10 54.2 49.2 46.7 46.6 57.2 52.5 52.4 49.7
15 55.5 51.0 47.9 46.9 59.6 55.0 53.7 51.9
20 54.7 50.5 48.4 47.1 57.8 53.4 52.3 50.5
Table 5: Ablation of the number of MANTA blocks on Breakfast.
Method Params. (M) Stages Blocks Total Blocks Inf. Time (sec)
GTDA [55] 3.9 5 9 45 71.8
Transf. (15) [5] 1.2 1 15 15 15.6
Transf. (18) [5] 1.4 1 18 18 18.2
Ours 1.4 1 15 15 1.1
Attention. To better understand the benefits of the proposed model, particularly the state-space model-based MANTA block, we explored the alternative of utilizing the attention-based Transformer (Transf.) [51] block instead. Like SSM-based models, transformers have a global temporal receptive field, but unlike SSMs, each element in the sequence has direct access to all others, avoiding the need for information compression at the cost of quadratic complexity. To mitigate the high complexity, window-based transformers [5, 31] limit their computation to a constrained set of values, making the computation sparse and more efficient.
For our experiments, we made use of such a Transformer designed to balance efficiency and performance for long-term dense temporal action segmentation. Specifically, we use the block from [5], which combines local windowed attention to ensure neighbouring consistency with global sparse ‘atrous’ attention to model the sequence’s long-term context. We compare our MANTA block to [5] in Tab. 7, replacing all blocks in our network with theirs. We evaluate two Transformer models with different total block counts to ensure a fair comparison. As shown in Tab. 6, while Transf.(15) with $15$ blocks matches the block count in the MANTA architecture, Transf.(18) with $18$ blocks matches the parameter count of our proposed network.
While both tested Transformer-based networks outperform the previously proposed GTDA in effectiveness and efficiency, among themselves, Transf.(18) is more effective than Transf.(15), but due to the larger number of layers, it suffers in efficiency. However, both Transformer models fall short compared to our proposed MANTA model. Specifically, our model is 16.5×16.5$\times$16.5 × faster than Transf.(18) and achieves, on average, $1.1\%$ and $2.4\%$ better Mean and Top-1 MoC.
Long-Range Sequence Modelling. In Fig. 5, we further examine how our proposed MANTA model performs in comparison to the other tested models across different video lengths. Specifically, we calculate the Average Mean MoC for videos grouped by their duration, i.e., we first compute the Mean MoC for each video individually and then average the results within each length-based group. While MANTA consistently outperforms both Transf.(18) and GTDA on all video lengths, the performance gap increases for longer videos. As discussed in Sec. 3, GTDA encounters difficulties with longer sequences due to the lack of global context caused by the sparse receptive fields of its layers. Despite maintaining a global receptive field across layers, Transf.(18) also yields inferior results compared to our proposed model. We hypothesize that the issue stems from the lack of strong local bias in the Transformer’s temporal updates, which contrasts with the MANTA blocks. The lack of such bias would require the network to learn it from the data, potentially requiring specifically tailored optimization procedures and/or larger training sets.

### v2

@@ 4.2 Protocol and Evaluation
Table 1: Comparison of MANTA to state-of-the-art methods on Breakfast, Assembly101 and 50Salads.
@@ 4.3 Results
Comparison to State of the Art. In this section, we present the results of our proposed model, comparing its performance with previous methods, as shown in Tab. 1. Our model achieves state-of-the-art results across all three datasets for both reported metrics. Specifically, on the Breakfast dataset, our method significantly outperforms previous approaches across all observation and anticipation horizons. On the Assembly101 dataset, MANTA surpasses all prior methods, with the exception of Top-1 MoC accuracy for $\beta=0.1$, where it performs slightly worse than GTDA. Lastly, on the 50Salads dataset, our approach sets a new state of the art for Mean MoC accuracy and Top-1 MoC accuracy for $\alpha=0.3$. For $\alpha=0.2$, our model’s Top-1 MoC is slightly lower compared to earlier methods. Generally, the average improvement in Top-1 MoC of MANTA compared to previous works across different anticipation ratios is consistently larger for $\alpha{=}0.3$ than for $\alpha{=}0.2$ for all datasets, indicating superior performance of our model for longer input sequences.
Beyond its effectiveness, our proposed model demonstrates substantial efficiency improvements compared to [57]. Fig. 4 compares the inference and training time of MANTA and the previous best-performing method GTDA [57] on the Breakfast dataset. On the left, it shows the average inference time required to generate $S=25$ samples with observation and anticipation ratios set to $\alpha=0.3$ and $\beta=0.5$ for videos in the test splits. While GTDA requires an average of $71.8$ seconds per video, our model needs only $1.1$ seconds, achieving a 65.3×65.3$\times$65.3 × speedup. Besides faster inference, our model significantly reduces training time as well, as shown in Fig. 4 on the right. The average runtime for a single epoch of GTDA is 379 seconds, while MANTA only takes 57 seconds per epoch. Additionally, due to the reduced number of the required blocks, our proposed model has 2.8×2.8$\times$2.8 × fewer parameters compared to GTDA, as shown in Tab. 6 and is more memory-efficient.
Finally, we show a qualitative comparison of our model to GTDA in Fig. 3. In this example, GTDA anticipates future actions wrong, while our proposed MANTA model predicts plausible action sequences. Here, the plausibility is judged by the number of predicted segments and their corresponding classes. GTDA produces a high number of segments and includes actions from unrelated recipes (e.g., pour cereal and break egg). This over-segmentation results from GTDA’s global but sparse receptive field of its later layers. In contrast, MANTA’s dense receptive field ensures consistency in the predicted segments and their classes. More qualitative results are in the supp. mat.
Figure 3: Qualitative comparison of MANTA and GTDA on Breakfast. Best viewed zoomed in.
@@ 4.4 Ablations
Firstly, the selectivity of the BSS layer is enabled by employing the S6 layer for temporal modelling, which applies input-specific SSM matrices for data-driven sequence processing. To assess the impact of selectivity, we replaced the S6 layer with S4, for which the parameters remain fixed across timesteps. The results in Tab. 2 show a significant performance drop without selectivity (Ours w/o Sel.). This is expected, as the input sequences to our model consist of two distinct segments: observed and masked future frames. Processing these segments homogeneously is inherently sub-optimal, since it is not possible to adjust the contribution of individual values regardless of their validity and/or importance. GTDA made similar observations, where selectivity was ensured by data-driven gating.
@@
MANTA Block. We further performed a series of experiments to explore how changes to the MANTA Block beyond the BSSL affect the performance of the final model. Firstly, we investigated the effect of the internal structure of the block, as shown in Tab. 4. Following recent works [17, 45], we experimented with modifying the channel mixing (FF) component of our block. Removing the mixing MLP altogether (Ours w/o FF) hinders the performance, as expected, due to a lack of channel-wise communication. However, introducing channel attention (Ours w/ SE), as proposed in [17], fails to improve the results and leads to a performance drop. Moreover, we experimented with introducing a dilated convolution layer with an exponentially growing dilation rate (Ours w/ Dil. Conv) before the LN layer, akin to [57]. This, however, led to the degradation of results. Finally, we evaluated the variant [26], which adds a residual block with normalization and MLP, as in a transformer block, after each block with BSSL. While it achieves better Top-1 MoC, Mean MoC is lower.
@@
Method Params. (M) Mem. (GB) Stages Blocks Total Blocks Inf. Time (sec)
GTDA [57] 3.9 19.2 5 9 45 71.8
Transf. (15) [5] 1.2 11.3 1 15 15 15.6
Transf. (18) [5] 1.4 13.2 1 18 18 18.2
Ours 1.4 10.2 1 15 15 1.1
Attention. To better understand the benefits of the proposed model, particularly the state-space model-based MANTA block, we explored the alternative of utilizing the attention-based Transformer (Transf.) [52] block instead. Like SSM-based models, transformers have a global temporal receptive field, but unlike SSMs, each element in the sequence has direct access to all others, avoiding the need for information compression at the cost of quadratic complexity. To mitigate the high complexity, window-based transformers [5, 32] constrain their computation to a limited set of values, making the computation sparse and therefore more efficient.
For our experiments, we made use of such a Transformer designed to balance efficiency and performance for long-term dense temporal action segmentation. Specifically, we use the block from [5], which combines local windowed attention to ensure neighbouring consistency with global sparse ‘atrous’ attention to model the sequence’s long-term context. We compare our MANTA block to [5] in Tab. 5, replacing all blocks in our network with theirs. We evaluate two Transformer models with different total block counts to ensure a fair comparison. As shown in Tab. 6, while Transf.(15) with $15$ blocks matches the block count in the MANTA architecture, Transf.(18) with $18$ blocks matches the parameter count of our proposed network. While both tested Transformer-based networks outperform the previously proposed GTDA in effectiveness and efficiency, among themselves, Transf.(18) is more effective than Transf.(15), but due to the larger number of layers, it suffers in efficiency. However, both Transformer models fall short compared to our proposed MANTA model. Specifically, our model is 16.5×16.5$\times$16.5 × faster than Transf.(18) and achieves, on average, $1.1\%$ and $2.4\%$ better Mean and Top-1 MoC.
Long-Range Sequence Modelling. In Fig. 5, we further examine how our proposed MANTA model performs in comparison to the other tested models across different video lengths. Specifically, we calculate the Average Mean MoC for videos grouped by their duration, i.e., we first compute the Mean MoC for each video individually and then average the results within each length-based group. While MANTA consistently outperforms both Transf.(18) and GTDA on all video lengths, the performance gap increases for longer videos. More specifically, MANTA shows a $3.8\%$ improvement in average Mean MoC over GTDA on shorter sequences, and a $5.5\%$ improvement on longer sequences. As discussed in Sec. 3, GTDA encounters difficulties with longer sequences due to the lack of global context caused by the sparse receptive fields of its layers. Despite maintaining a global receptive field across layers, Transf.(18) also yields inferior results compared to our proposed model. We hypothesize that the issue stems from the lack of strong local bias in the Transformer’s temporal updates, which contrasts with the MANTA blocks. The lack of such bias would require the network to learn it from the data, potentially requiring specifically tailored optimization procedures and/or larger training sets.
@@ 5 Conclusion
Acknowledgments
The work has been supported by the ERC Consolidator Grant FORHUE (101044724), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) GA 1927/4-2 (FOR 2535 Anticipating Human Behavior), and the project iBehave (receiving funding from the programme “Netzwerke 2021”, an initiative of the Ministry of Culture and Science of the State of NorthrhineWestphalia). The authors gratefully acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy, through EuroHPC Regular Access Call- proposal No. EHPC-REG-2024R01-076. The authors also gratefully acknowledge the granted access to the Marvin cluster hosted by the University of Bonn. The sole responsibility for the content of this publication lies with the authors. We want to thank Mohamad Hakam Shams Eddin for his assistance with creating the teaser figure. The manta icons used in our paper were generated using Dall-E.
[…]
Supplementary Material
Here, we present additional dataset and implementation details, as well as additional quantitative and qualitative results for our proposed MANTA model. More precisely, we discuss the implementation of our model in Sec. 6 and provide additional details about the utilized datasets in Sec. 7. Next, in Sec. 8 we present additional ablation studies for MANTA. Finally, in Sec. 9 we present additional qualitative comparisons of MANTA to previous work.
6 Implementation details
We implemented our model using Pytorch. As per Tab. 9, we use a total of $B=15$ MANTA blocks for our final model. Our proposed network is trained for 90 epochs using the Adam [24] optimizer with a learning rate of $0.0005$ for Breakfast and Assembly, and $0.001$ for 50Salads. Following [57], we use $T=1000$ diffusion steps for training, $D=50$ DDIM steps for inference on Breakfast and Assembly101, and $D=10$ inference steps for 50Salads.
7 Datasets
In Tab. 7, we show additional details for the datasets used in our work. Specifically, we provide average and maximum video durations, as well as the average and maximum number of individual segments per video. Since in our adapted anticipation protocol only up to $50\%$ of the video frames are utilized for anticipation, we additionally provide the statistics for the corresponding intervals of the videos in brackets in blue, including only frames falling into the anticipation intervals. As one can observe, the long temporal horizon of videos used for future anticipation and the numerous action segments that need to be predicted highlight the long-term nature of the addressed task.
Dataset Avg. Num. Seg. Max. Num. Seg Avg. Dur. (min) Max. Dur. (min)
Breakfast 7 (3) 25 (15) 2.3 (1.2) 10.8 (5.4)
50Salads 20 (11) 26 (18) 6.4 (3.2) 10.1 (5.1)
Assembly101 12 (5) 73 (40) 3.5 (1.8) 25.0 (12.5)
Table 7: (Left) Number of segments and (Right) duration for the whole video and in the anticipation interval.
8 Ablation Study
8.1 Bidirectional State-Space Layer (BSSL)
In Tab. 2 and Tab. 3 of the main paper, we analyzed the selectivity and bi-directionality of the proposed BSS layer. Here, we examine how independent forward and backward scanning contributes to the final performance of the MANTA model. More specifically, we tested if having independent parameters for forward and backward scanning branches is the best way to structure the BSS layer. To investigate this, we evaluated the effect of weight-sharing between the two branches. As shown in Tab. 8, while Top-1 MoC accuracy is similar across networks with shared and independent weights, Mean MoC accuracy is higher in the model with branch-specific weights. We, therefore, keep the weights separate for the two BSSL branches.
MoC Shared $\beta\ (\alpha=0.2)$ $\beta\ (\alpha=0.3)$
0.1 0.2 0.3 0.5 0.1 0.2 0.3 0.5
Mean 28.5 23.4 23.8 22.9 33.2 27.7 28.2 27.1
27.7 25.3 24.6 23.8 34.2 30.9 29.1 27.7
Top-1 53.7 51.0 48.8 46.7 60.7 55.9 53.3 50.5
55.5 51.0 47.9 46.9 59.6 55.0 53.7 51.9
Table 8: Ablation of weight sharing for forward and backward branches of the BSSL on Breakfast.
8.2 MANTA Block
We experimented with varying the total number of blocks in the final model (Tab. 9). Empirically, we found that the model with $B=15$ blocks showed the best results, with further increase or decrease in the number of blocks harming the model’s performance.
MoC Num. $\beta\ (\alpha=0.2)$ $\beta\ (\alpha=0.3)$
blocks 0.1 0.2 0.3 0.5 0.1 0.2 0.3 0.5
Mean 10 26.6 24.3 23.5 23.1 32.7 29.4 28.1 26.8
15 27.7 25.3 24.6 23.8 34.2 30.9 29.1 27.7
20 27.2 24.5 23.4 23.2 32.1 29.3 27.7 26.6
Top-1 10 54.2 49.2 46.7 46.6 57.2 52.5 52.4 49.7
15 55.5 51.0 47.9 46.9 59.6 55.0 53.7 51.9
20 54.7 50.5 48.4 47.1 57.8 53.4 52.3 50.5
Table 9: Ablation of the number of MANTA blocks on Breakfast.
8.3 Samples.
We analyze the effect of the total sample count on MANTA’s performance in Tab. 10. The number of samples has a marginal effect on the Mean MoC accuracy whereas the Top-1 MoC increases with the number of samples. While this is expected, as Top-1 MoC only considers the sample that is closest to the ground-truth, the increase in Top-1 MoC with a higher number of samples demonstrates the diversity of the generated predictions. Otherwise, the Top-1 MoC would saturate after a small number of samples.
MoC Num. $\beta\ (\alpha=0.2)$ $\beta\ (\alpha=0.3)$
samples 0.1 0.2 0.3 0.5 0.1 0.2 0.3 0.5
Mean 5 27.7 25.6 24.7 23.9 33.9 30.7 28.9 27.5
15 27.7 25.4 24.6 23.8 34.2 30.9 29.0 27.8
25 25.5 25.3 24.6 23.8 34.2 30.9 29.1 27.7
Top-1 5 42.3 39.3 37.1 36.1 47.5 44.2 41.4 40.4
15 51.4 47.3 44.7 44.0 55.8 52.0 50.5 48.6
25 55.5 51.0 47.9 46.9 59.6 55.0 53.7 51.9
Table 10: Ablation of the number of samples on Breakfast.
8.4 Robustness
We report the standard deviation for the MANTA model, computed over 3 seeds on the Breakfast dataset, in Tab. 11. As expected, the std. is higher for Top-1 MoC, but the values are low, indicating the robustness of our proposed model.
MoC $\beta\ (\alpha=0.2)$ $\beta\ (\alpha=0.3)$
0.1 0.2 0.3 0.5 0.1 0.2 0.3 0.5
Mean 0.07 0.07 0.06 0.05 0.04 0.18 0.09 0.11
Top-1 0.5 0.6 0.3 0.2 0.3 0.6 0.8 0.5
Table 11: Standard deviation of MANTA on Breakfast dataset computed over 3 runs with different seeds.
9 Qualitative Results
We provide qualitative comparisons of our proposed MANTA model to the previous best-performing GTDA [57] on the Breakfast dataset in Figs. 6-8, on the 50Salads dataset in Fig. 9 and on Assembly101 dataset in Fig. 10.
Figure 6: Qualitative comparison of MANTA (top) to GTDA [57] (bottom) on the Breakfast dataset. Best viewed zoomed in.
Figure 7: Qualitative comparison of MANTA (top) to GTDA [57] (bottom) on the Breakfast dataset. Best viewed zoomed in.
Figure 8: Qualitative comparison of MANTA (top) to GTDA [57] (bottom) on the Breakfast dataset. Best viewed zoomed in.
Figure 9: Qualitative comparison of MANTA (top) to GTDA [57] (bottom) on the 50Salads dataset. Best viewed zoomed in.
Figure 10: Qualitative comparison of MANTA (top) to GTDA [57] (bottom) on the Assembly101 dataset. Best viewed zoomed in.